# V2P: From Background Suppression to Center Peaking for Robust GUI Grounding Task

### v1

@@ 3.4 Valley-to-Peak Training
$\mathcal{L}=\mathcal{L}_{\text{NTP}}+\lambda_{1}\mathcal{L}_{\text{Sup\_Attn}}+\lambda_{2}\mathcal{L}_{\text{Action\_Attn}},$ (12)
@@ 4 Experiment
4.1 Experiment Setup
Implementation Details.
We use Qwen2.5-VL-7B-Instruct Bai et al. (2025) as the backbone VLM. Training is conducted on 32 NVIDIA H200 GPUs for 1 epoch. The key hyperparameters are a learning rate of 5e-6, a global batch size of 128, and $\sigma=1$ for the Fitts-Gaussian Peak Modeling.
Our data recipe is built from several public GUI datasets, the source data totaling approximately 1M screenshots. To ensure annotation quality, we apply Ominiparser Wan et al. (2024) to detect bounding boxes for all samples and filter those where the IoU between ground truth (GT) and parser-detected boxes is less than 0.3, as such cases likely contain annotation errors, this step improves the data consistency for training. More details about the implementation are provided in Appendix A.
Evaluation Benchmarks & Metric.
We conduct experiments to assess the performance of our methods on two widely adopted GUI visual grounding benchmarks: ScreenSpot-v2 Wu et al. (2024b), and ScreenSpot-Pro Li et al. (2025). Among these, ScreenSpot-Pro stands out due to its focus on higher-resolution user interfaces and increased domain variability, such as industrial applications and multi-window setups. This dataset serves as a valuable resource for evaluating generalization capabilities in realistic scenarios. To quantify performance, we utilize the Element Accuracy metric, which calculates the percentage of cases where the predicted click point resides within the target element’s ground-truth bounding box. Further details about these benchmarks are available in Appendix B
4.2 Main Result
Our proposed V2P-7B achieves outstanding performance on both the ScreenSpot-v2 and ScreenSpot-Pro benchmarks, delivering significant improvements over existing approaches. As shown in Tab. 1, V2P-7B achieves an average accuracy of 50.54% on the challenging ScreenSpot-Pro dataset, outperforming all GUI-specific models, including RL-based methods such as SE-GUI-7B (47.3%) and GUI-$\text{G}^{2}$-7B (47.5%). Notably, our model secures the highest scores in 6 out of 12 task categories, with exceptional performance in icon-based interactions (25.83% average), where precise spatial grounding is crucial. On ScreenSpot-v2 (Tab. 2), V2P-7B achieves an impressive average accuracy of 92.3%, with remarkable performance in Mobile-Text (98.1%) and Web-Text (95.4%) scenarios.
These advancements are driven by V2P’s dual optimization strategy: (1) the Suppression Attention mechanism effectively mitigates background distractions, while (2) Fitts-Gaussian Labeling addresses center-edge confusion, significantly improving localization in icon-based tasks. Remarkably, V2P obtains these results using supervised fine-tuning alone, avoiding the complexity of reinforcement learning pipelines employed by recent RL-based models. This streamlined training process not only simplifies implementation but also leaves substantial room for future performance enhancements through potential RL integration. The consistent gains across diverse UI platforms (mobile, desktop, web) and interaction types (text, icon) affirm V2P’s robust generalizability for real-world GUI grounding applications.

### v3

@@ 3.4 Valley-to-Peak Training
$\mathcal{L}=\mathcal{L}_{\text{NTP}}+\lambda_{1}\mathcal{L}_{\text{Sup\_Attn}}+\lambda_{2}\mathcal{L}_{\text{Action\_Attn}}$ (14)
@@ 4 Experiment
4.1 Experimental Setup
We utilize Qwen2.5-VL-Instruct (both 7B and 3B) (Bai et al., 2025) as backbones. To ensure a rigorously fair comparison and isolate algorithmic contributions, we strictly follow the data recipe of the baseline GUI-Actor (Wu et al., 2025), with a learning rate of 5e-6 and $\sigma=1.0$. Comprehensive details are provided in App. A.
We evaluate on a comprehensive suite of six benchmarks. Our primary evaluation focuses on ScreenSpot-v2 (Wu et al., 2024) and ScreenSpot-Pro (Li et al., 2025), as they provide the most standardized assessment across diverse platforms and challenging high-resolution OOD scenarios.
To further verify robustness and agentic potential, we also test on OSWorld-G (Xie et al., 2025), UI-Vision (Element Grounding) (Nayak et al., 2025), UI-I2E (Liu et al., 2025), and MMBench-GUI L2 (Liu et al., 2024).
4.2 Main Results
Tab. 1 presents a comprehensive evaluation of V2P against other baselines.
Superior Performance on ScreenSpot Benchmarks. As our primary evaluation field, V2P-7B demonstrates exceptional capabilities among models of similar scale. On ScreenSpot-v2, it achieves a competitive accuracy of 92.4%. More critically, on the high-difficulty ScreenSpot-Pro, which features high-resolution screens and OOD applications, V2P-7B attains 52.5%, significantly outperforming the strong baseline GUI-Actor-7B (44.6%) and UI-TARS-72B (38.1%). This substantial margin validates that V2P’s attention calibration is particularly effective in handling the dense, visually complex interfaces typical of professional GUI environments.
Generalization to Agentic Scenarios. To assess the model’s potential as a perception backend for autonomous agents, we extend our evaluation to four benchmarks featuring interaction traces and functional reasoning requirements: OSWorld-G (Xie et al., 2025), UI-Vision (Element Grounding) (Nayak et al., 2025), UI-I2E (Liu et al., 2025), and MMBench-GUI L2 (Liu et al., 2024). As shown in Tab. 1, V2P-7B demonstrates superior performance across the majority of evaluations. Notably, V2P-7B surpasses all other baselines on UI-Vision, UI-I2E, and MMBench-GUI L2. This consistent superiority highlights the model’s exceptional functional reasoning and semantic understanding. Furthermore, on OSWorld-G, V2P matches the specialist JEDI-7B (Xie et al., 2025) (52.5%) despite using only $\sim$50k PC samples versus JEDI’s millions. Moreover, V2P significantly surpasses JEDI on other benchmarks, highlighting superior data efficiency and generalization beyond specific domains.
Scalability and Efficiency. As shown in the Controlled Comparison group, V2P-3B consistently outperforms its direct competitor GUI-Actor-3B across all six benchmarks. Notably, on some challenging benchmarks, it even surpasses significantly larger scale models. This result underscores the pure algorithmic superiority of the V2P framework and its consistent effectiveness across varying model scales.
Model General Grounding Complex & Semantic Grounding
ScreenSpot-v2 ScreenSpot-Pro OSWorld-G UI-Vision UI-I2E MMBench
Proprietary & General VLMs
GPT-4o 80.7 0.8 1.38 2.87
Operator 70.5 40.6
Qwen2.5-VL-3B 80.9 16.1 27.3 41.7
Qwen2.5-VL-7B 88.8 26.8 31.4 0.85 53.8 33.9
GUI-Specialized Models (SFT)
SeeClick-9.6B 55.1 1.1 5.39 26.4
OS-Atlas-7B 84.1 18.9 27.7 9.02 58.6 41.4
Aguvis-7B 86.0 22.9 38.7 13.7 53.2 45.7
UGround-V1-7B 87.6 31.1 36.4 12.9 70.3 65.7
UI-TARS-7B 91.6 35.7 47.5 17.6 61.4 64.3
JEDI-7B 91.7 39.5 54.1 24.8
UI-TARS-72B 90.3 38.1 57.1 25.5 73.7 74.3
Controlled Comparison (Identical Training Data)
GUI-Actor-3B 91.0 42.2 45.9 21.9 63.7 73.5
V2P-3B (Ours) 91.4 48.5 48.8 26.0 69.5 77.6
GUI-Actor-7B 92.1 44.6 49.3 24.3 68.2 76.5
V2P-7B (Ours) 92.4 52.5 52.5 28.8 75.6 79.9
Table 1: Main Results Comparison. We evaluate V2P against state-of-the-art baselines across six diverse benchmarks, covering general, high-resolution, and agentic GUI scenarios. V2P-7B significantly outperforming baselines under comparable settings.
4.3 Ablation and Analysis
4.3.1 Component Ablation Study
To validate the necessity of our proposed modules, we conducted a standard ablation study on ScreenSpot-Pro (Tab. 2). Removing Fitts-Gaussian Peak Modeling (FGPM) leads to a significant performance drop of 5.0%, confirming its critical role in precise localization. Further removing Suppression Attention (SA) results in an additional loss of 3.2%. These results verify that both modules are indispensable for the V2P framework.
Model Variant Pro Avg. $\Delta$
V2P-7B (Full) 52.5 -
w/o FGPM 47.5 -5.0
w/o FGPM & SA 44.3 -8.2
Table 2: Component Ablation on ScreenSpot-Pro. Both FGPM and SA contribute significantly to the final performance.
4.3.2 Attribution of Performance Gains
To investigate the underlying reasons for V2P’s superior performance, we conducted a quantitative performance gains attribution analysis on 182 samples where V2P-7B successfully corrected the failures of the baseline GUI-Actor-7B (Wu et al., 2025). As shown in Tab. 3, the results reveal that 50.5% of the performance gains stem from effectively suppressing Background Distraction, while 35.7% are attributed to resolving Center-Edge Confusion. This provides strong empirical evidence that V2P’s dual-loss mechanism functions exactly as designed.
Baseline Error Type Count Contribution
Background Distraction 92 50.5%
Center-Edge Confusion 65 35.7%
Other / Normal Attention 25 13.7%
Total Improved Samples 182 100%
Table 3: Performance Gains Attribution Analysis. We analyzed samples from ScreenSpot-Pro where V2P-7B made correct predictions while the baseline GUI-Actor (Wu et al., 2025) failed. The majority of gains come from correcting background and center-edge errors.
Figure 3: Gaussian Factor and Generalization Ability Analysis. (a) Impact of Gaussian Factor $\sigma$. A smaller $\sigma$ (sharper peak) benefits precision, with the optimal performance achieved at $\sigma=1.0$ for ScreenSpot-Pro. Larger $\sigma$ values degrade performance due to introduced label noise. (b, c) Generalization Ability. V2P shows consistent improvement, whereas the baseline suffers from overfitting on OOD data.
4.3.3 Performance Leap on Tiny Targets
To evaluate performance on fine-grained targets, we categorized UI elements across ScreenSpot-v2 and ScreenSpot-Pro based on their area relative to the patch size $n$ ($14\times 14$). Specifically, elements are classified as Small ($n\leq A<4n$), Medium ($4n\leq A<9n$), and Large ($A\geq 9n$). As shown in Tab. 4, V2P-7B outperforms the baseline GUI-Actor-7B (Wu et al., 2025) by 10.0% on these small elements. This demonstrates the superiority of V2P in the fine-grained positioning of small targets.
Furthermore, the data shown in Tab. 4 also reveals a critical distribution shift between benchmarks: ScreenSpot-v2 is dominated by large elements (size >9n>9n), which offer vast spatial tolerance. Consequently, even spatially diffuse attention maps often fall within these generous boundaries, which explains the high accuracy of the baseline on ScreenSpot-v2, effectively masking its inherent localization imprecision. In contrast, ScreenSpot-Pro is densely populated with small elements that tolerate negligible error. Consequently, V2P-7B’s precision advantage, while masked on the coarse-grained ScreenSpot-v2, is fully realized on the challenging ScreenSpot-Pro.
4.3.4 Sensitivity to Gaussian Factor $\sigma$
To analyze the impact of the Gaussian factor $\sigma$ on grounding precision, we conducted ablation experiments on ScreenSpot-v2 and ScreenSpot-Pro across varying $\sigma$ values. As shown in Fig. 3(a), model performance is strongly sensitive to this hyperparameter. On ScreenSpot-v2, accuracy improves from 91.3% ($\sigma=6.0$) to 92.4% ($\sigma=0.5$). Similarly, ScreenSpot-Pro achieves its peak accuracy of 52.5% at $\sigma=1.0$, while larger $\sigma$ values cause a significant decline.
Element Size ScreenSpot-v2 ScreenSpot-Pro
GUI-Actor V2P GUI-Actor V2P
Small ($n\sim 4n$) 50.0% 60.0% 17.5% 23.8%
Medium ($4n\sim 9n$) 71.4% 85.7% 43.1% 47.9%
Large (>9n>9n) 93.2% 92.9% 60.3% 66.6%
Table 4: Size-stratified Performance. V2P achieves substantial gains on small elements in ScreenSpot-v2 and ScreenSpot-Pro, underscoring its superior capability in precise fine-grained localization.
We attribute this phenomenon to the spatial concentration of the attention mechanism. Larger $\sigma$ values generate broader Gaussian distributions, which tend to dilute the spatial focus and introduce background noise into the attention maps. Conversely, a smaller $\sigma$ produces sharper Gaussian peaks. This acts as a tight spatial constraint, allowing the model to localize UI elements with higher precision and resulting in more accurate click predictions. These results underscore the necessity of balancing $\sigma$: while excessively large values hinder localization, a moderately small $\sigma$ (e.g., 1.0) significantly enhances spatial accuracy.
4.3.5 Training Stability and Generalization
Finally, we evaluate the training stability of V2P-7B compared to the Aguvis-7B (Xu et al., 2025). As visualized in Fig. 3(b) and (c), V2P-7B demonstrates a consistently ascending accuracy curve on both in-distribution (ScreenSpot-v2) and out-of-distribution (ScreenSpot-Pro) benchmarks. In sharp contrast, Aguvis-7B (Xu et al., 2025) exhibits a distinct "overfitting-to-distribution" pattern: while its performance improves on ScreenSpot-v2, it suffers from a continuous performance decline on the OOD ScreenSpot-Pro after the 20% training milestone. This confirms that our human-like visual attention mechanism (Fitts-Gaussian Peak Modeling and Suppression Attention) effectively mitigates the overfitting inherent to textual coordinate supervision, ensuring robust generalization across unseen scenarios.
5 Conclusion
In this paper, we address the critical bottlenecks of Background Distraction and Center-Edge Confusion in GUI grounding by proposing Valley-to-Peak (V2P) framework. Mimicking human visual processing, V2P synergizes Suppression Attention to eliminate background noise and Fitts-Gaussian Peak Modeling to construct sharp, size-adaptive peaks at actionable centers.
By emulating this human-like strategy for visual localization, our approach fosters a more authentic spatial understanding of complex interfaces. Extensive experiments confirm the effectiveness of this framework: V2P achieves exceptional results on ScreenSpot-v2 (92.4%) and the challenging ScreenSpot-Pro (52.5%), consistently outperforming existing strong baselines. Notably, our method demonstrates remarkable robustness on fine-grained small targets and out-of-distribution scenarios, effectively bridging the gap between coarse perception and precise actuation. By enabling agents to "see" and "focus" like human users, V2P offers a scalable and robust foundation for the next generation of general-purpose GUI agents.
Limitations
While V2P demonstrates exceptional performance across various benchmarks, several limitations remain to be addressed:
  • Ambiguity among Semantically Similar Targets: As analyzed in our failure case studies (see App. D), the model occasionally struggles when multiple UI elements share high semantic similarity, such as identical icons with different functional purposes. This suggests that visual calibration alone may not fully resolve deep logical intent without more comprehensive UI context.
  • Generalization to Unconventional Designs: The model’s attention distribution can become highly dispersed when encountering unconventional or cluttered layouts that deviate from the training distribution, indicating uncertainty in complex visual environments.
  • Computational Overhead: The introduction of the self-attention module to enhance spatial coherence among visual patches may introduce marginal increases in inference latency compared to simple coordinate regression methods, particularly when processing high-resolution screenshots with a large number of patches.
Ethics Statement
In this work, we propose the Valley-to-Peak (V2P) framework to improve GUI grounding by mimicking human visual processing. We adhere to the ACL Code of Ethics and highlight the following:
  • Data Privacy: All training and evaluation datasets used in this study are from publicly available academic sources. We have strictly followed data recipe guidelines to exclude samples containing personal identifiable information (PII).
  • Mitigation of Bias: Our training data spans multiple operating systems and platforms (Mobile, Desktop, Web) to minimize algorithmic bias toward specific UI design patterns.
Acknowledgements
The authors would like to thank the anonymous reviewers for their insightful feedback. We acknowledge the use of generative AI tools for polishing the linguistic quality and refining the prose of this manuscript. All technical claims and final content remain the sole responsibility of the authors.
Appendix A Training and Inference Details
A.1 Source Training Data
Following GUI-Actor (Wu et al., 2025), we compile our training dataset from several publicly available, high-quality GUI datasets, with summary statistics provided in Tab. 5. To ensure fair evaluation, we also exclude any samples from Wave-UI that overlap with the test sets of downstream tasks.
Dataset # of Elements # of Screenshots Platform
Uground Web–Hybrid (Gou et al., 2025) 8M 775K Web
GUI-Env (Chen et al., 2025) 262K 70K Web
GUI-Act (Chen et al., 2025) 42K 13K Web
AndroidControl (Li et al., 2024) 47K 47K Android
AMEX (Chai et al., 2025) 1.2M 100K Android
Wave-UI 50K 7K Hybrid
Total 9.6M 1M
Table 5: Overview of training datasets used for GUI-Actor.
Appendix B Benchmarks
Our evaluation centers on six sophisticated benchmarks for GUI visual grounding:
ScreenSpot-v2 (Wu et al., 2024) encompasses 1,272 carefully annotated instructions, each paired with corresponding target elements across diverse GUI environments, including mobile (Android and iOS), desktop (macOS and Windows), and web platforms. The dataset is designed to improve the quality and reliability of GUI visual grounding tasks, addressing key challenges such as eliminating ambiguities in natural language instructions and resolving annotation errors. By refining the alignment between textual descriptions and interface elements, ScreenSpot-v2 provides a robust and standardized benchmark for evaluating grounding models.
ScreenSpot-Pro (Li et al., 2025), meanwhile, focuses on more demanding scenarios, especially those involving high-resolution professional applications. It contains 1,581 tasks annotated by domain experts across 23 specialized software applications, spanning three operating systems. This benchmark significantly broadens the scope of GUI visual grounding by introducing interfaces with industrial software and multi-window layouts, creating a larger domain gap compared to most pretraining data. With its increased complexity and domain diversity, ScreenSpot-Pro is an invaluable resource for assessing the generalization ability of models in realistic and challenging GUI environments.
OSWorld-G is the grounding-specific subset derived from the OSWorld benchmark (Xie et al., 2025), a unified evaluation environment for multimodal agents on Ubuntu. Unlike static datasets, OSWorld-G consists of screenshots captured from a fully functional, interactive operating system. It evaluates the model’s ability to localize actionable elements within dynamic and complex real-world desktop workflows, serving as a direct proxy for an agent’s practical utility in autonomous computer control tasks.
UI-Vision (Element Grounding) (Nayak et al., 2025) is designed to rigorously test the semantic understanding of user interface elements. While standard grounding tasks often rely on text matching (OCR), UI-Vision focuses on functional icons and visual symbols (e.g., identifying a "magnifying glass" as "search" or a "floppy disk" as "save") that lack explicit textual labels. Performance on this benchmark reflects the model’s capacity for visual reasoning and its ability to interpret the functional affordances of GUI components.
UI-I2E (Image-to-Element) (Liu et al., 2025) evaluates the capability to parse the hierarchical structure of a screen. The task requires the model to map raw pixel inputs to structured representations, effectively "reading" the underlying layout or accessibility tree of the interface. High accuracy on UI-I2E indicates that the model possesses a deep understanding of UI composition and element spatial relationships, rather than merely memorizing surface-level patterns.
MMBench-GUI L2 (Liu et al., 2024) is the GUI-specific subset (L-2 category) of the massive MMBench suite. Adopting a robust CircularEval strategy with multiple-choice questions, it assesses fine-grained perception and reasoning abilities within graphical interfaces. This benchmark serves as a standardized indicator of the model’s general-purpose multimodal intelligence in the GUI domain, complementing the pure localization metrics of ScreenSpot.
Model ScreenSpot-v2 Accuracy (%)
Proprietary Models
Operator 47.3 41.5 90.2 80.3 92.8 84.3 70.5
GPT-4o + OmniParser-v2 95.5 74.6 92.3 60.9 88.0 59.6 80.7
General Open-source Models
Qwen2.5-VL-3B 93.4 73.5 88.1 58.6 88.0 71.4 80.9
Qwen2.5-VL-7B 97.6 87.2 90.2 74.2 93.2 81.3 88.8
GUI-specific Models (SFT)
SeeClick-9.6B 78.4 50.7 70.1 29.3 55.2 32.5 55.1
Magma-8B 62.8 53.4 80.0 57.9 67.5 47.3 61.5
OS-Atlas-4B 87.2 59.7 72.7 46.4 85.9 63.1 71.9
UI-TARS-2B 95.2 79.1 90.7 68.6 87.2 78.3 84.7
OS-Atlas-7B 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Aguvis-7B 95.5 77.3 95.4 77.9 91.0 72.4 86.0
UGround-V1-7B 95.0 83.3 95.0 77.8 92.1 77.2 87.6
UI-TARS-72B 94.8 86.3 91.2 87.9 91.5 87.7 90.3
GUI-Actor-3B 97.6 83.4 96.9 83.6 94.0 85.7 91.0
UI-TARS-7B 96.9 89.1 95.4 85.0 93.6 85.2 91.6
GUI-Actor-7B 97.6 88.2 96.9 85.7 93.2 86.7 92.1
GUI-specific Models (RL)
SE-GUI-7B - - - - - - 90.3
LPO-8B - - - - - - 90.5
Ours
V2P-7B 98.1 88.0 96.1 89.7 95.4 84.4 92.4
Appendix C Detailed Experimental Results on ScreenSpot-v2 and ScreenSpot-Pro
We provide extended experimental results, including fine-grained performance breakdowns and comparisons against a broader set of baselines. Detailed statistics are presented in Tab. 6 and Tab. 7.